OpenStudy (anonymous):

you are shopping for school clothes. You want to buy 2 shirts that cost $10.40 each. use the distributive property to mentally calculate the cost of the shirts.

OpenStudy (anonymous):

2x = 10.4 x = 5.2

OpenStudy (anonymous):

noo.

OpenStudy (anonymous):

whyyyyy?

OpenStudy (anonymous):

Total = n*(unitary) Total=2*10.4=20.8